### v1

# Modelling subsonic turbulence with SPH-EXA

Abstract
The numerical simulation of subsonic turbulence with SPH has traditionally been hindered by E0 errors, inaccurate gradient evaluations, and excessive dissipation. In this work, we present numerical results of SPH simulations of subsonic turbulence and compare to state-of-the-art codes such as AREPO and GIZMO. For the first time, SPH can reproduce the results of such codes, showing a similar interval of wavenumber in the inertial range of the Kolmogorov cascade in the subsonic regime. We use the SPH-EXA code to perform these simulations, a GPU-based state-of-the-art SPH code with high performance and extreme scalability at its core. SPH-EXA is coupled with a modern SPH solver based on methods such as an integral approach to gradient calculation, artificial viscosity switches that include a linear field cleaner, a flexible family of pairing-resistant interpolation kernels, generalized volume elements, and a controlled handling of density jumps which maximizes Lagrangian compatibility. In addition, it includes a novel and extremely scalable gravity solver for astrophysical applications.
I Introduction
Smoothed Particle Hydrodynamics (SPH) is a very flexible and adaptable numerical technique for solving partial differential equations in continuum mechanics that has been successfully applied in many different fields. In particular, its application in fluid dynamics, solid mechanics, and astrophysics has substantially benefited those respective fields since its conception around year 1980 [1, 2]. At the core of its success lay its excellent conservation properties, the lack of an underlying structured mesh, and its natural treatment of three-dimensional calculations [3, 4, 5]. Despite these considerable advantages, SPH has historically struggled with certain types of calculation. In particular, those related to the development of subsonic turbulence.
Turbulence is a fundamental multiscale phenomenon in fluid dynamics, characterized by chaotic and unpredictable fluid flow [6]. It occurs in both natural environments and technological applications, making its study and understanding crucial across various fields of science and engineering. Turbulence plays a vital role in the formation and evolution of astronomical structures, from accretion disks around black holes to the large-scale structure of the universe. It influences the distribution of galaxies, star formation, and the dynamics of the interstellar medium. In meteorology, turbulence affects weather patterns, cloud formation, and the dispersion of pollutants. Understanding turbulence is essential for accurate weather forecasting, climate modeling, and assessing environmental impacts. Ocean currents exhibit turbulent behavior, affecting marine ecosystems, climate systems (through heat and carbon exchange), and navigation. Turbulence also plays a role in phenomena such as El Niño and the mixing of nutrients from deep waters to the surface. In aerospace, automotive, and chemical engineering, controlling or exploiting turbulence can lead to more efficient designs. For example, reducing turbulence around aircraft wings or vehicle bodies can reduce fuel consumption and emissions. In contrast, enhancing turbulence in reactors can improve mixing and chemical reactions. Turbulence challenges our understanding of nonlinear dynamical systems and fluid mechanics. It is a rich field for theoretical and computational research, seeking universal laws and models to describe turbulent flows. Despite its ubiquity and importance, turbulence remains one of the most challenging areas of physics to model and predict due to its complex, non-linear nature. Progress in this field beyond the heuristic but very successful pioneering model by Kolmogorov [7] can lead to advances in technology, improved environmental management, and deeper insights into the fundamental processes of nature.
The numerical simulation of turbulence has been a very powerful tool to progress in its understanding, yet such simulations are very challenging as they require to be accurate across a vast range of scales for very long simulation times. This implies high-resolution simulations with state-of-the-art codes. In general, there are no significant discrepancies between simulations with SPH and with grid-based methods for supersonic turbulence [8]. Nevertheless, reproducing subsonic or transonic turbulence with SPH has been traditionally a challenge because it failed to produce results with sufficient quality to compete with grid-based methods for a similar resolution [9]. First studies of subsonic turbulence with SPH in the context of cosmological simulations [10, 11] which found a rough agreement with the Kolmogorov power spectra but only in a very limited range of wavenumbers. More recent attempts to simulate turbulence by direct numerical simulation (DNS) only led to minor [9] or moderate improvements [12], with the latter highlighting the importance of reducing the viscosity away from shocks. However, the velocity power spectra still showed significant differences with those obtained with mesh-based calculations. All these studies identified zeroth order, $E0$, errors, excessive dissipation, and inaccurate estimation of gradients as the principal causes of the difficulties in simulating turbulence with SPH at low-Mach numbers.
Adapting Large Eddy Simulation (LES) concepts to SPH was shown to be a promising route to simulating turbulence [13, 14] although at the price of introducing additional, and somehow artificial, terms in the basic SPH formulation with a loss in the Lagrangian compatibility of the method. On another note, there is little doubt that DNS simulations of subsonic turbulence would benefit from a better estimation of gradients, [15], e.g. by using an integral approach (IA) to derivatives [16] to obtain a better agreement with the Kolmogorov power spectra and in a wider range of wavelengths.
In this work, we present SPH simulations of subsonic turbulence calculated with the SPH-EXA code, a state-of-the-art, highly scalable hydrodynamic code that incorporates the latest improvements in High-performance Computing (HPC), such as full GPU porting of the SPH computation kernels, efficient domain decomposition, GPU-based neighbor search, on-the-fly initial conditions generation, and in-situ visualization. It also incorporates the latest improvements in the SPH technique such as the mentioned IA to calculate gradients, generalized volume elements, optimized Lagrangian compatibility [17, 18] and an efficient linear field cleaner [19] among others. Equipped with such a bag of resources we show that the simulations of turbulence with SPH-EXA can compete with those obtained with top reference codes such as GIZMO [20] and AREPO [21]. Our results reproduce an increasing inertial range of the Kolmogorov cascade with resolution, which, thanks to the excellent scaling properties of SPH-EXA, will make an increasingly larger fraction of the inertial range of turbulence accessible to DNS in the forthcoming years.
II Methodology
II-A Gradients
SPH-EXA implements an integral approach to calculate derivatives [16, 22].
The IA formalism calculates gradients of an arbitrary function $f$ by solving the following matrix equation:
$\left[\begin{array}[]{c}\partial f/\partial x_{1}\\ \partial f/\partial x_{2}\\ \partial f/\partial x_{3}\\ \end{array}\right]_{a}=\left[\begin{array}[]{ccc}\tau_{11}&\tau_{12}&\tau_{13}\\ \tau_{21}&\tau_{22}&\tau_{23}\\ \tau_{31}&\tau_{32}&\tau_{33}\end{array}\right]^{-1}_{a}\left[\begin{array}[]{c}I_{1}\\ I_{2}\\ I_{3}\\ \end{array}\right]_{a}\,,$ (1)
where the $\tau$ tensor elements are defined as:
$\tau_{ij,a}=\sum_{b}V_{b}~{}(x_{i,b}-x_{i,a})(x_{j,b}-x_{j,a})~{}W_{ab}(h_{a});~{}i,j=1,3\,.$ (2)
Here $x_{i}$ is the vector position coordinates, $V_{b}$ the volume element, $W_{ab}$ the interpolation kernel, and $I_{i,a}$ the vector integral elements, which are calculated as:
$I_{i,a}=\left[\sum_{b}V_{b}f_{b}~{}(x_{i,b}-x_{i,a})~{}W_{ab}(h_{a})\right];~{}i=1,3\,.$ (3)
As a consequence, the gradients calculation is:
$\nabla_{i}f_{a}=\sum_{b}V_{b}(f_{b}-f_{a})A_{i,ab}(h_{a});~{}i=1,3\,$ (4)
where,
$A_{i,ab}(h_{a})=\sum_{j=1}^{3}c_{ij,a}(h_{a})(x_{j,b}-x_{j,a})W_{ab}(h_{a})\,$ (5)
where the summation is up to 3 for 3D simulations and $c_{ij,a}$ the elements of the inverse of matrix $\mathcal{T}$, that is, $\mathcal{C}=\mathcal{T}^{-1}$ (Eq. 1).
II-B Artificial Viscosity
The basic AV implementation in SPH-EXA is the following:
$\Pi_{ab}^{AV}=\begin{cases}-\frac{\alpha}{2}v_{ab}^{sig}~{}w_{ab}&\text{for~{}~{}$\mathbf{x}_{ab}\cdot\mathbf{v}_{ab}<0$}\\ 0&\text{otherwise}\end{cases}\,,$ (6)
where $\alpha$ is the viscous parameter and is usually set to 1. The signal velocity estimator between a pair of particles is:
$v^{sig}_{ab}=\frac{1}{2}(c_{a}+c_{b})-2w_{ab}\,,$ (7)
Finally, $c_{a}$ and $c_{b}$ are the local speed of sound for each particle, and $w_{ab}=(\mathbf{v}_{ab}\cdot\mathbf{x}_{ab})/|\mathbf{x}_{ab}|$.
A linear velocity field implies a smooth flow without the presence of shocks. Therefore, linear velocity fields should have a vanishing AV. We reconstruct the local velocity field for each particle by removing the linear component of the velocity jump that is used to trigger and calculate the AV in Eq. 6 [19]. Assuming that it is differentiable in the neighborhood of each SPH particle, the best linear approximation of the velocity field is given by multiplying its Jacobian by the distance to its neighbors. Instead of using $\mathbf{v}_{ab}=\mathbf{v}_{a}-\mathbf{v}_{b}$ to calculating $w_{ab}$, we use $\mathbf{v}^{\prime}_{ab}=\mathbf{v}^{\prime}_{a}-\mathbf{v}^{\prime}_{b}$, defined as:
$v^{\prime}_{a,i}$ $\equiv v_{a,i}-\frac{1}{2}\phi_{ab}\mathbf{J}_{\mathbf{v}_{a}}\mathbf{x}^{T}_{ab}$ (8)
$v^{\prime}_{b,i}$ $\equiv v_{b,i}+\frac{1}{2}\phi_{ba}\mathbf{J}_{\mathbf{v}_{b}}\mathbf{x}^{T}_{ab}$ (9)
where $\mathbf{x}^{T}_{ab}=\mathbf{x}_{a}-\mathbf{x}_{b}$ is the column vector of relative distance and $\mathbf{J}_{\mathbf{v}}$ is the Jacobian of the velocity field. In the case of the presence of discontinuities, such a linear estimator is not accurate enough; hence, the $\phi_{ab}$ factor is designed to go to zero in these cases, while it is kept close to 1 when the field is smooth. To this end, we use a van Leer-like limiter [19]:
$\phi_{ab}$ $=\max\left[0,\min\left(1,\frac{4F_{ab}}{(1+F_{ab})^{2}}\right)\right]\times\kappa_{ab}\,,$ (10)
$\kappa_{ab}$$\equiv\begin{dcases}\exp{\left[-\left(\frac{q_{ab}-q_{crit}}{q_{fold}}\right)^{2}\right]}&\text{if~{}~{}$q_{ab}<q_{crit}$}\,,\\ 1&\text{if~{}~{}$q_{ab}\geq q_{crit}$}\,,\end{dcases}$ (11)
$q_{ab}$ $\equiv\min(q_{a},q_{b})\,,$ (12)
$q_{crit}$ $\equiv\left(\frac{32\pi}{3nb_{a}}\right)^{1/3}\,,$ (13)
$F_{ab}$ $\equiv\frac{\mathbf{x}_{ab}\mathbf{J}_{\mathbf{v}_{a}}\mathbf{x}^{T}_{ab}}{\mathbf{x}_{ab}\mathbf{J}_{\mathbf{v}_{b}}\mathbf{x}^{T}_{ab}}\,,$ (14)
where $q_{a}$ is the normalized distance between neighbors ($q_{a}=|\mathbf{x}_{b}-\mathbf{x}_{a}|/h_{a}$), $q_{crit}$ is the average of the interparticle distance normalized to the smoothing length in 3D, and $nb_{a}$ is the number of neighbors of particle $a$. The parameter $q_{fold}$ determines the width of the correction factor $\kappa_{ab}$ (Eq. 11) that is applied to the van Leer limiter (Eq. 10) and we use the recommended value [19] of $q_{fold}=0.2$.
The final step is to effectively control excessive dissipation. This can be done by allowing a dynamic individual viscous parameter $\alpha_{a}$ in Eq. 6, rather than having a constant value for all SPH particles. Several methods have been developed to calculate $\alpha_{a}$, but the nominal works of [23] and [24] are the inspiration behind our implementation in SPH-EXA. The main idea is that $\alpha_{a}$ should be initialized by default at a rather low value (usually $\alpha_{a}=0.05$). Then it increases rapidly if dissipation is needed and decays exponentially to its original value when it is not needed, effectively acting as a switch for the AV at the position of each SPH particle.
We start by calculating a local viscous parameter $\alpha_{loc,a}$, following the definition [23]:
$\alpha_{loc,a}=\begin{dcases}\frac{\alpha_{max}\mathcal{D}_{a}}{\mathcal{D}_{a}+h_{a}|\nabla\cdot\mathbf{v}_{a}|+0.05c_{a}}&\text{for~{}~{}$\nabla\cdot\mathbf{v}_{a}<0$}\,,\\ 0&\text{otherwise}\,,\end{dcases}$ (15)
where $\mathcal{D}_{a}=h^{2}_{a}|\nabla(\nabla\cdot\mathbf{v}_{a})|$ predicts flow convergence by comparing the gradient of $\nabla\cdot\mathbf{v}_{a}$ with the local value of $\nabla\cdot\mathbf{v}_{a}$. If the former is large enough and the fluid is actually converging ($\nabla\cdot\mathbf{v}_{a}<0$), $\alpha_{loc,a}$ adopts a value between 0 and $\alpha_{max}$. Furthermore, the term $0.05c_{a}$, where $c_{a}$ stands for the local speed of sound, acts as a base value for the size of the velocity fluctuations that can trigger dissipation, and it also works as a term to prevent divergent $\alpha_{loc,a}$.
Once we know $\alpha_{loc,a}$ we have two options:
  • if $\alpha_{loc,a}\geq\alpha_{a}$: we instantaneously rise the dissipation setting $\alpha_{a}=\alpha_{loc,a}$ ,
  • if $\alpha_{loc,a}<\alpha_{a}$: we let $\alpha_{a}$ decay smoothly .
In the case that $\alpha_{a}$ has to decay, we evaluate an instantaneous decay rate $\dot{\alpha}_{a}$ as in [23]:
$\dot{\alpha}_{a}=\begin{dcases}(\alpha_{loc,a}-\alpha_{a})/\tau_{a}&\text{if~{}~{}$\alpha_{min}<\alpha_{loc,a}<\alpha_{a}$}\,,\\ (\alpha_{min}-\alpha_{a})/\tau_{a}&\text{if~{}~{}$\alpha_{min}\geq\alpha_{loc,a}$}\,.\end{dcases}$ (16)
The decay time $\tau_{a}$ is defined as $\tau_{a}=h_{a}/v_{sig,max}$, and it is a function of the local spatial resolution and the maximum signal velocity $v_{sig,max}$ in the neighborhood of particle $a$, calculated according to Eq. 7. The effect of Eq. 16 is to allow alpha to decay exponentially to $\alpha_{loc,a}$, or $\alpha_{min}$ if $\alpha_{loc,a}$ is too low. In this way, we ensure a minimum dissipation that helps reduce the random noise that might arise from particle disorder. However, we note that even though all particles have at least a minimum value $\alpha_{min}$, dissipation is automatically cut off when $\mathbf{r}_{ab}\cdot\mathbf{v}_{ab}\geq 0$, as stated in Eq. 6. Also, there is no restriction for $\alpha_{min}$, which can be set to zero.
II-C Volume elements
In the context of SPH, where we are discretizing the SPH integral interpolant, it becomes necessary to compute the volume element that corresponds to each particle. An uncomplicated yet highly effective method is $dx^{\prime 3}\equiv V_{b}\sim m_{b}/\rho_{b}$. However, it is known [25, 26] that the discretization of the volume element, present in most SPH equations, can be improved beyond this classical approximation using generalized volume elements (GVE).
GVE are usually expressed as:
$V_{a}=\frac{X_{a}}{\sum_{b=1}^{nb}X_{b}W_{ab}}\,,$ (17)
where $X$ is an arbitrary function. In SPH-EXA we adopted the GVE described in the work of [18], which has shown to provide a better treatment of discontinuities, stronger resistance to tensile instability, better partition of unity, and more accurate gradient evaluation:
$X_{a}=\frac{m_{a}}{\rho^{0}_{a}}\,,$ (18)
where $\rho^{0}_{a}=\sum_{b}m_{b}W_{ab}$ is the standard SPH density calculation. With Eqs. 17 and 18 we can simply calculate the particle density as $\rho_{a}=m_{a}/V_{a}$.
II-D Lagrangian compatibility
One of the most common artifacts in SPH simulations is tensile instability, which can, in some cases, prevent mixing and the development of hydrodynamical instabilities. Several approaches have been developed to deal with such problems [27, 28, 25], all of which require a departure to some extent from the pure Lagrangian formulation. In SPH-EXA we adopted a novel implementation [18], which, to our knowledge, is currently the only adaptive method that sacrifices Lagrangian compatibility exclusively in regions of the fluid at (and adjacent to) contact interphases, improving the accuracy at those regions, in turn. Hence, we define a self-adaptive parameter that is individual to each particle-neighbor pair as follows:
$\sigma_{ab}\equiv\begin{cases}0&\text{if~{}~{}$At_{ab}\leq At_{min}$}\,,\\ R(At_{ab}-At_{min})&\text{if~{}~{}$At_{min}\leq At_{ab}\leq At_{max}$}\,,\\ 1&\text{if~{}~{}$At_{ab}\geq At_{max}$}\,,\end{cases}$ (19)
with $R=(At_{max}-At_{min})^{-1}$ and $At_{ab}$ is the pairwise Atwood number, defined as $At_{ab}=\left|\frac{\rho_{a}-\rho_{b}}{\rho_{a}+\rho_{b}}\right|$. We use the values $At_{max}=0.2$ and $At_{min}=0.1$, as suggested in [18]. In this implementation $0\leq\sigma\leq 1$ (see Eqs. 22 and 23). In particular, when $\sigma=0$ the resulting SPH equations are fully Lagrangian, while $\sigma=1$ leads to a version less Lagrangian compatible but more tensile resistant. The parameter $\sigma$ appears in the momentum and energy SPH equations and has the effect of crossing and uncrossing the particle and neighbor’s magnitudes.
II-E Interpolation kernel
We use the sinc family ($S_{n}$) [29, 30] as interpolation kernels,
$W_{ab}=S_{n}=\frac{K}{h_{a}^{d}}\begin{dcases}1&q_{a}=0\\ \left[\operatorname{sinc}\left(\frac{\pi}{2}q_{a}\right)\right]^{n}&q_{a}\leq 2\\ 0&q_{a}>2\\ \end{dcases}$ (20)
where $K$ is a normalization constant, $h$ is the smoothing length, $d$ is the spatial dimension, $q_{a}$ is the normalized distance between neighbor particles ($q_{a}=|\mathbf{x}_{b}-\mathbf{x}_{a}|/h_{a}$), and $n$ is a user-defined real value that makes the kernel more centrally peaked as it increases. Usually $93<n<$ 9. It is known [31] that a necessary condition for an SPH kernel to avoid pairing instability is that its Fourier transform is definite positive. As such, they proposed the use of Wendland functions [32] as SPH kernels. These have been adopted by many researchers for this reason. The other side of the coin is that Wendland kernels need a large number of neighbors to perform accurate interpolations, which leads to considerable computational efforts. A possible alternative is to use pairing-resistant kernels. Despite not having an always positive Fourier transform, they are constructed to have small negative Fourier regions and push them to large wavenumbers. This allows these kernels to use a lower number of neighbors than Wendland functions, while still suppressing pairing instability. Sinc kernels are an example of such kernels. The simulations reported in this work were calculated choosing $n=6$ in Eq. 20.
II-F SPH-EXA hydrodynamic equations
The final set of hydrodynamic equations in SPH-EXA is as follows:
$\rho_{a}=$ $m_{a}/V_{a}\,,$ (21)
$\begin{split}\frac{dv_{i,a}}{dt}=&-\sum_{b}m_{b}\times\\ &\left[\frac{X_{a}^{2-\sigma}X_{b}^{\sigma}P_{a}}{\Omega_{a}m_{a}^{2}~{}k_{a}}A_{i,ab}(h_{a})+\frac{X_{b}^{2-\sigma}X_{a}^{\sigma}P_{b}}{\Omega_{b}m_{b}^{2}~{}k_{b}}A_{i,ab}(h_{b})\right]\,,\end{split}$ (22)
$\begin{split}\frac{du_{a}}{dt}=&\frac{X_{a}^{2-\sigma}P_{a}}{m_{a}^{2}\Omega_{a}k_{a}}\sum_{b}\sum^{d}_{i=1}m_{b}X_{b}^{\sigma}\left[(\mathbf{v}_{a}-\mathbf{v}_{b})\cdot\mathbf{A}_{ab}(h_{a})\right]\\ &+\left(\frac{du_{a}}{dt}\right)^{AV}\,.\end{split}$ (23)
Here $P_{a}$ is the pressure, $X_{a}=m_{a}/\rho^{0}_{a}$ the selected function for the GVE, $k_{a}$ the normalization of the GVE ($k_{a}\equiv\sum_{b}X_{b}W_{ab}$ in Eq. 17), $A_{i,ab}(h_{a})$ the IA term, and $\Omega_{a}$ the grad-h term. The parameter $\sigma$ controls how close the implementation is to a pure Lagrangian scheme (as discussed above). The last term in the energy equation includes the contributions of artificial viscosity.
II-G Initial conditions
We have designed an embedded initial conditions (IC) generator in SPH-EXA. In all our tests, we use a template of a relaxed box with $50^{3}$ particles. This box has previously been relaxed to achieve a glass-like configuration [33] with constant density and periodic boundary conditions. Replicating this box in different directions allows us to create on-the-fly initial conditions with a user-defined number of particles and in a distributed manner, only limited by the memory of the system. In this way, ICs can be scaled up just by changing one input argument and without the need to store large files, ensuring reproducibility and simplifying the process of performing suites of tests at scale.
All simulations presented here were performed on a box of side $L=1$, periodic boundary conditions, initial density $\rho=1$, and ’quasi’-isothermal equation of state with an adiabatic index $\gamma=1.001$. The fluid is stirred using the Ornstein-Uhlenbeck process [34, 35] following a parabolic amplitude field $\sim 1-(k-k_{c})^{2}$, with $k=|\mathbf{k}|$, and $k_{c}=2$, which corresponds to a length $L/2$, and vanishing for $k\leq 1$ and $k\geq 3$. This approach is implemented to reduce the impact of the stirring on the energy cascade, facilitating a natural transfer of energy from larger to smaller scales. We also use a natural mixing of $1/3$ compressive and $2/3$ solenoidal contributions. All simulations start in a homogenous state and are stirred for a total of $t=10$, until they reach the stationary state [^1] All our times are normalized with the sound-crossing time, which in our setup is $t_{sc}=L/c=1$. with an RMS Mach number of $M=0.3$.
III Results
To evaluate the weight and significance of the different elements that contribute to our SPH implementation, we conducted identical simulations while progressively incorporating more of these elements. Figure 1 shows the power spectrum of the velocity field of simulations of $250^{3}$ particles with increasing complexity, compared to the S3 simulation of [9] with $256^{3}$ particles and the classic SPH formulation (i.e. gradients calculated directly as kernel derivatives, no GVEs, and no viscosity switches), labeled STD in our plot.
Figure 1: Comparison of the power spectra of the velocity field for $M=0.3$ with different elements incorporated into the SPH solver. The dashed line is the theoretical Kolmogorov power spectrum of $\sim k^{-5/3}$. STD stands for standard SPH and is taken from [9] (simulation S3 in their Fig. 6, with $256^{3}$ particles). All other simulations were performed with SPH-EXA and $250^{3}$ particles. IA stands for including the integral formalism to calculate derivatives. VAV stands for variable AV and implies the inclusion of AV switches. GVE is for generalized volume elements, and AVC for AV cleaning. All these elements have been described in Sec. II.
The inclusion of IA recovers the correct slope in the inertial range of the Kolmogorov cascade for smaller wavenumbers, considerably extending it compared to STD, but is not enough to affect the classical excessive slope of $\sim-4$ at large wavenumbers. At these small scales something else is needed. The inclusion of GVE and artificial viscosity switches (labeled as VAV in Fig. 1 provides a step in the right direction. With these elements, we start to see the bottleneck effect as the excess of energy at increasing wavenumber, commonly encountered close to the resolution limit in mesh-based simulations. Finally, the inclusion of the artificial viscosity cleaner (AVC in Fig. 1) provides the final push, extending the inertial range almost a decade with respect STD, showing the bottleneck effect clearly, and having a slope in the dissipative regime similar to those of non-Lagrangian codes (see Fig. 4).
Figure 2 presents the power spectrum of the velocity field for simulations performed with SPH-EXA and with different particle counts at the end of each simulation ($t=10$). It is clear that with increasing resolution, the inertial range over the Kolmogorov scale is extended to larger wavenumber, and more importantly, that the rate of enhancement with increasing number of particles is not marginal or small, but considerable and consistent with that obtained with modern mesh-based and moving-mesh methods. This is a novel result, which means that SPH can indeed resolve smaller scales with higher resolution, in opposition to the results found with older SPH implementations by [9]. Figure 3 shows a thin slice of the velocity field for increasing resolution with SPH-EXA, from $200^{3}$ up to $3000^{3}$, where it is evident how we can resolve a delicate and rich mixture of large- and small-scale features.
Figure 2: Comparison of the power spectra of the velocity field at $t=10$ for $M=0.3$ at increasing resolution. Each line corresponds to an increase of a factor 2 in resolution with respect to the previous curve.
Figure 3: Slice of the module of the velocity field with increasing resolution and calculated with SPH-EXA.
Now we can directly compare our results with other state-of-the-art codes that are not purely Lagrangian. Figure 4 displays a comparison of the power spectra calculated with AREPO, GIZMO, and SPH-EXA. As can be seen, SPH-EXA is able to reproduce the same inertial range. We can also see the bottleneck effect to a very similar extent in all three codes. As this artificial excess of energy happens at a very similar wavenumber for similar resolutions, it points to the fact that the transition from the inertial range to the dissipation range happens at the same scale for a given resolution and for all three codes. It is worth noting that in Fig. 4 we compare codes with a similar amount of fluid elements, which does not necessarily correspond to same resolution. In SPH codes, the spatial resolution is defined by the smoothing length, which covers significantly more than just one particle. At $256^{3}$ cells, the spatial resolution of AREPO roughly corresponds to $1/256\sim 3.9\times 10^{-3}$, which is in fact similar to our $400^{3}$ SPH simulation, which has an average smoothing length of $h\sim 3.6\times 10^{-3}$. This means that if we compare at the same resolution, and according to the results of Fig. 4, SPH-EXA captures a considerably larger inertial range of the Kolmogorov cascade than AREPO. Likewise, the simulation of AREPO with $512^{3}$ cells has an approximate resolution of $\sim 1.9\times 10^{-3}$, which corresponds to our $800^{3}$ simulation ($h\sim 1.8\times 10^{-3}$) represented with a dashed line in Fig. 4.
Figure 4: Comparison of the power spectra of the velocity field for $M=0.3$ for different codes. AREPO [9] and GIZMO [20] curves are from their respective papers.
Figure 5 shows, from top to bottom, the module of the velocity, density, and enstrophy of the fluid at the end of our simulations. The two columns on the left are an excerpt of the results with AREPO and the SPH used in [9]. It is clear from these two columns that the SPH used in that work was not able to resolve subsonic turbulence, since it is missing all small-scale features. The two columns on the right show the results of SPH-EXA with $250^{3}$ and $1000^{3}$ particles. The similarities with the results of AREPO are evident, and there is a rich mix of fine structures that becomes richer with increasing resolution. This can also be seen in Fig. 3, which illustrates a slice of the module of the velocity field at $t=10$ for different resolutions.
Figure 5: Comparison of SPH-EXA results with AREPO and SPH-tav in [9]. From top to bottom: module of velocity, density, and enstrophy. The two columns on the left are an excerpt of the original results of AREPO and their SPH with reduced artificial viscosity (both with $256^{3}$ fluid elements) of [9] shown in their Fig. 4. The two columns on the right are the results of SPH-EXA with $250^{3}$ and $1000^{3}$
Finally, Fig. 6 is a detailed image of a slice of the module of velocity of the $3000^{3}$ simulation where we can appreciate the very small-scale features resolved with SPH-EXA.
Figure 6: Slice of the module of the velocity field at $t=10$ for a simulation with $3000^{3}$ particles. Color follows the same scale as Fig. 3 but in grey.
IV Summary and conclusions
SPH has matured enough to be competitive in simulations of subsonic turbulence, obtaining results comparable to Eulerian codes. In particular, SPH-EXA combines a modern SPH formulation with the latest techniques in parallelization and the use of accelerators. This allowed us to overcome historical limitations of SPH, such as excessive dissipation and inaccurate gradient evaluations, which have traditionally prevented SPH from matching the performance of grid-based and moving-mesh codes in this field. Our findings not only align well with results from non-Lagrangian codes like AREPO and GIZMO, but also illustrate the robustness of SPH in handling complex turbulent flows at resolutions previously thought unattainable with this method.
We demonstrated that incorporating various components that address problems at multiple simulation scales collectively achieves these results. Isolated enhancements in the SPH framework cannot tackle the subsonic turbulence; however, a synergy of these improvements can. In particular, our combination of an integral approach to derivatives, generalized volume elements, artificial viscosity switches, and artificial viscosity cleaner shows excellent results.
Finally, we proved that the inertial range of the power spectra significantly spans over larger vawenumbers with increasing resolution, exhibiting the same features as other non-Lagrangian state-of-the-art codes. These improvements have allowed SPH to finally tap into its inherent advantages, such as its intrinsic Lagrangian nature, excellent conservation properties, and adaptability to varied physical scenarios, thus paving the way for its application in more demanding astrophysical contexts.
Acknowledgment
This work has been supported by the Swiss Platform for Advanced Scientific Computing (PASC) project ”SPH-EXA: Optimizing Smoothed Particle Hydrodynamics for Exascale Computing”. It has also been carried out as part of the SKACH consortium through funding from SERI and supported by the Spanish MINECO grant PID2020-117252GB-100 and by the AGAUR/Generalitat de Catalunya grant SGR-386/2021. We also acknowledge the financial support of the European Research Council through the ERC Synergy Grant “ECOGAL” (project ID 855130), from the German Excellence Strategy via the Heidelberg Cluster of Excellence (EXC 2181 - 390900948) “STRUCTURES”, and from the German Ministry for Economic Affairs and Climate Action in project “MAINN” (funding ID 50OO2206). The numerical calculations have been supported by the EuroHPC-JU Extreme Scale Access Mode through the project ”TGSF: The role of turbulence and gravity in star formation: unveiling the sonic scale with Smoothed Particle Hydrodynamics” (proposal ID: EHPC-EXT-2023E01-031). The authors acknowledge the support of the LUMI Supercomputer Center, the Swiss National Supercomputing Center (CSCS - allocation c32), and the Center for Scientific Computing (sciCORE) at the University of Basel, where these calculations were performed.

### v2

@@
-Modelling subsonic turbulence with SPH-EXA
-Abstract
-The numerical simulation of subsonic turbulence with SPH has traditionally been hindered by E0 errors, inaccurate gradient evaluations, and excessive dissipation.
-In this work, we present numerical results of SPH simulations of subsonic turbulence and compare to state-of-the-art codes such as AREPO and GIZMO.
-For the first time, SPH can reproduce the results of such codes, showing a similar interval of wavenumber in the inertial range of the Kolmogorov cascade in the subsonic regime.
-We use the SPH-EXA code to perform these simulations, a GPU-based state-of-the-art SPH code with high performance and extreme scalability at its core.
-SPH-EXA is coupled with a modern SPH solver based on methods such as an integral approach to gradient calculation, artificial viscosity switches that include a linear field cleaner, a flexible family of pairing-resistant interpolation kernels, generalized volume elements, and a controlled handling of density jumps which maximizes Lagrangian compatibility.
-In addition, it includes a novel and extremely scalable gravity solver for astrophysical applications.
-I Introduction
-Smoothed Particle Hydrodynamics (SPH) is a very flexible and adaptable numerical technique for solving partial differential equations in continuum mechanics that has been successfully applied in many different fields.
-In particular, its application in fluid dynamics, solid mechanics, and astrophysics has substantially benefited those respective fields since its conception around year 1980 [1, 2].
-At the core of its success lay its excellent conservation properties, the lack of an underlying structured mesh, and its natural treatment of three-dimensional calculations [3, 4, 5].
-Despite these considerable advantages, SPH has historically struggled with certain types of calculation.
-In particular, those related to the development of subsonic turbulence.
-Turbulence is a fundamental multiscale phenomenon in fluid dynamics, characterized by chaotic and unpredictable fluid flow [6].
-It occurs in both natural environments and technological applications, making its study and understanding crucial across various fields of science and engineering.
-Turbulence plays a vital role in the formation and evolution of astronomical structures, from accretion disks around black holes to the large-scale structure of the universe.
-It influences the distribution of galaxies, star formation, and the dynamics of the interstellar medium.
-In meteorology, turbulence affects weather patterns, cloud formation, and the dispersion of pollutants.
-Understanding turbulence is essential for accurate weather forecasting, climate modeling, and assessing environmental impacts.
-Ocean currents exhibit turbulent behavior, affecting marine ecosystems, climate systems (through heat and carbon exchange), and navigation.
-Turbulence also plays a role in phenomena such as El Niño and the mixing of nutrients from deep waters to the surface.
-In aerospace, automotive, and chemical engineering, controlling or exploiting turbulence can lead to more efficient designs.
-For example, reducing turbulence around aircraft wings or vehicle bodies can reduce fuel consumption and emissions.
-In contrast, enhancing turbulence in reactors can improve mixing and chemical reactions.
-Turbulence challenges our understanding of nonlinear dynamical systems and fluid mechanics.
-It is a rich field for theoretical and computational research, seeking universal laws and models to describe turbulent flows.
-Despite its ubiquity and importance, turbulence remains one of the most challenging areas of physics to model and predict due to its complex, non-linear nature.
-Progress in this field beyond the heuristic but very successful pioneering model by Kolmogorov [7] can lead to advances in technology, improved environmental management, and deeper insights into the fundamental processes of nature.
-The numerical simulation of turbulence has been a very powerful tool to progress in its understanding, yet such simulations are very challenging as they require to be accurate across a vast range of scales for very long simulation times.
-This implies high-resolution simulations with state-of-the-art codes.
-In general, there are no significant discrepancies between simulations with SPH and with grid-based methods for supersonic turbulence [8].
-Nevertheless, reproducing subsonic or transonic turbulence with SPH has been traditionally a challenge because it failed to produce results with sufficient quality to compete with grid-based methods for a similar resolution [9].
-First studies of subsonic turbulence with SPH in the context of cosmological simulations [10, 11] which found a rough agreement with the Kolmogorov power spectra but only in a very limited range of wavenumbers.
-More recent attempts to simulate turbulence by direct numerical simulation (DNS) only led to minor [9] or moderate improvements [12], with the latter highlighting the importance of reducing the viscosity away from shocks.
+
+11institutetext: Center for Scientific Computing - sciCORE, University of Basel, Klingelberstrasse 61, 4056 Basel, Switzerland
+11email: ruben.cabezon@unibas.ch
+22institutetext: Center for Data Analytics - CeDA, University of Basel, Klingelberstrasse 61, 4056 Basel, Switzerland 33institutetext: Departament de Física.
+Universitat Politècnica de Catalunya (UPC).
+Av.
+Eduard Maristany 16, 08019 Barcelona, Spain
+33email: domingo.garcia@senior.upc.edu
+44institutetext: Institut d’Estudis Espacials de Catalunya (IEEC), 08860 Castelldefels (Barcelona), Spain 55institutetext: Universität Heidelberg, Zentrum für Astronomie, Institut für Theoretische Astrophysik, Albert-Ueberle-Str.
+2, 69120 Heidelberg, Germany 66institutetext: Department of Mathematics and Computer Science, University of Basel, Spiegelgasse 1, 4051 Basel, Switzerland 77institutetext: Swiss National Supercomputing Center (CSCS), Via Trevano 131, 6900 Lugano, Switzerland 88institutetext: Department of Astrophysics, University of Zürich, Winterthurerstrasse 190, 8057 Zürich, Switzerland 99institutetext: Universität Heidelberg, Interdisziplinäres Zentrum für Wissenschaftliches Rechnen, Im Neuenheimer Feld 225, 69120 Heidelberg, Germany
+From particles to precision
+Simulating subsonic turbulence with Smoothed Particle Hydrodynamics
+Rubén M. Cabezón    Domingo García-Senz    Oliver Avril    Osman Seckin Simsek    Sebastian Keller    Axel S. Lechuga    Lucio Mayer    Ralf Klessen    Florina M. Ciorba
+(Received September 30, 20XX)
+Direct numerical simulation of subsonic turbulence with smoothed particle hydrodynamics (SPH) has traditionally been hampered by zeroth-order (E0) errors, inaccurate gradient evaluations, and excessive numerical dissipation.
+We demonstrate that the recently developed SPH-EXA code can overcome these challenges, producing results similar to those of other state-of-the-art non-SPH codes, such as AREPO and GIZMO.
+We show that SPH-EXA accurately reproduces the Kolmogorov inertial range scaling in the subsonic regime with increasing resolution, closely matching state-of-the-art non-SPH methods.
+This is an outstanding result given the important role of SPH codes in astrophysics and cosmology, and the ubiquity of turbulent phenomena.
+Key Words.:
+hydrodynamics – turbulence – Methods: numerical
+[…]
+1 Introduction
+Smoothed Particle Hydrodynamics (SPH) is a mesh-free, conservative method for continuum mechanics widely used in many different fields since its conception around 1980 (Lucy, 1977; Gingold and Monaghan, 1977).
+Its strengths include excellent conservation properties and a natural treatment of fully three-dimensional flows (Liu and Liu, 2010; Price, 2012b).
+Despite these advantages, SPH has historically struggled in some regimes, most notably in accurately simulating subsonic turbulence.
+Turbulence is a fundamental multiscale phenomenon in fluid dynamics, characterized by chaotic and nonlinear motion (Frisch, 1995).
+It occurs in both natural environments and technological applications, making its study central across science and engineering.
+In astrophysics, turbulence influences processes ranging from accretion flows to the interstellar medium, shaping star formation, galaxy evolution, and the emergence of large-scale structure.
+Despite its ubiquity and importance, turbulence remains one of the most challenging phenomena to model and predict because of its complex, non-linear nature.
+Progress in this field beyond the heuristic but very successful pioneering model of Kolmogorov (Kolmogorov, 1941) can lead to advances in technology, improved environmental management, and deeper insight into fundamental processes of nature.
+The direct numerical simulation (DNS) of turbulence requires high-resolution and state-of-the-art codes.
+In general, there are no significant discrepancies between SPH simulations and grid-based methods for supersonic turbulence (Kitsionas et al., 2009).
+Nevertheless, reproducing subsonic or transonic turbulence with SPH has traditionally been a challenge because it failed to produce results with sufficient quality to compete with grid-based methods for a similar resolution (Bauer and Springel, 2012).
+The first studies of subsonic turbulence with SPH in the context of cosmological simulations (Dolag et al., 2005; Valdarnini, 2011) found rough agreement with the Kolmogorov power spectra but only in a very limited range of wavenumbers.
+More recent DNS attempts led to minor (Bauer and Springel, 2012) or moderate improvements (Price, 2012a), with the latter highlighting the importance of reducing the viscosity away from shocks after showing that the Reynolds number is proportional to the Mach number in SPH simulations.
 However, the velocity power spectra still showed significant differences with those obtained with mesh-based calculations.
-All these studies identified zeroth order, $E0$, errors, excessive dissipation, and inaccurate estimation of gradients as the principal causes of the difficulties in simulating turbulence with SPH at low-Mach numbers.
-Adapting Large Eddy Simulation (LES) concepts to SPH was shown to be a promising route to simulating turbulence [13, 14] although at the price of introducing additional, and somehow artificial, terms in the basic SPH formulation with a loss in the Lagrangian compatibility of the method.
-On another note, there is little doubt that DNS simulations of subsonic turbulence would benefit from a better estimation of gradients, [15], e.g.
-by using an integral approach (IA) to derivatives [16] to obtain a better agreement with the Kolmogorov power spectra and in a wider range of wavelengths.
-In this work, we present SPH simulations of subsonic turbulence calculated with the SPH-EXA code, a state-of-the-art, highly scalable hydrodynamic code that incorporates the latest improvements in High-performance Computing (HPC), such as full GPU porting of the SPH computation kernels, efficient domain decomposition, GPU-based neighbor search, on-the-fly initial conditions generation, and in-situ visualization.
-It also incorporates the latest improvements in the SPH technique such as the mentioned IA to calculate gradients, generalized volume elements, optimized Lagrangian compatibility [17, 18] and an efficient linear field cleaner [19] among others.
-Equipped with such a bag of resources we show that the simulations of turbulence with SPH-EXA can compete with those obtained with top reference codes such as GIZMO [20] and AREPO [21].
-Our results reproduce an increasing inertial range of the Kolmogorov cascade with resolution, which, thanks to the excellent scaling properties of SPH-EXA, will make an increasingly larger fraction of the inertial range of turbulence accessible to DNS in the forthcoming years.
-II Methodology
-II-A Gradients
-SPH-EXA implements an integral approach to calculate derivatives [16, 22].
+All of these studies identified zeroth-order errors ($E0$), excessive dissipation, and inaccurate gradient estimation as the main causes of the difficulties in simulating turbulence with SPH in low-Mach numbers.
+In this work, we present SPH simulations of subsonic turbulence with SPH-EXA [^1] https://github.com/sphexa-org/sphexa
+https://deepwiki.com/sphexa-org/sphexa
+(Cavelan et al., 2020; Keller et al., 2023), a GPU-native state-of-the-art and highly scalable hydrodynamic code that incorporates the latest advances in high performance Computing (HPC), such as efficient domain decomposition, GPU-based neighbor search, on-the-fly initial conditions generation, and in-situ visualization.
+It also incorporates the latest improvements in the SPH technique, such as an integral approach to calculate gradients, generalized volume elements, and a slope-limited reconstruction (SLR) of the velocity field to control dissipation, among other improvements (see Appendices A and B).
+Equipped with such a set of resources, we show that the simulations of subsonic turbulence with SPH-EXA can compete with those obtained with top reference codes such as AREPO (Springel, 2010) and GIZMO (Hopkins, 2015).
+Our results show that, as resolution increases, we recover a growing portion of the inertial range of the Kolmogorov cascade.
+Owing to the excellent scaling properties of SPH-EXA, this will allow DNS to probe an ever larger fraction of the turbulent inertial range in the coming years.
+Figure 1: Comparison of the power spectra of the velocity field with different elements incorporated into the SPH solver.
+The dashed line is the theoretical Kolmogorov power spectrum of $\sim k^{-5/3}$.
+STD denotes standard SPH and is taken from (Bauer and Springel, 2012, simulation S3 in their Fig.
+6, with $256^{3}$ particles).
+All other simulations were performed with SPH-EXA and $250^{3}$ particles.
+The meaning of the remaining acronyms is provided in Sect.
+3.
+Figure 2: Comparison of the instantaneous power spectra of the velocity field at $t=10$ at increasing resolution.
+Each line corresponds to an increase of a factor of 2 in resolution with respect to the previous curve.
+2 Initial conditions
+We have designed an embedded initial conditions (IC) generator in SPH-EXA.
+In all our tests, we use a template of a relaxed box with $50^{3}$ particles.
+This box has previously been relaxed to achieve a glass-like configuration (Arth et al., 2019) with constant density and periodic boundary conditions.
+Replicating this box in different directions allows us to create on-the-fly initial conditions with a user-defined number of particles and in a distributed manner, only limited by the memory of the system.
+All simulations use a periodic box of side $L=1$ with uniform initial density $\rho=1$ and pressure $P=1$, and ’quasi’-isothermal equation of state ($\gamma=1.001$).
+Turbulence is driven by stochastic forcing modes at different wavelengths k in the range $1<|\mathbf{k}|<3$, with amplitudes following a parabolic distribution $\propto 1-(k-k_{c})^{2}$ peaking at $k_{c}=2$, ensuring that kinetic energy is injected only at the largest scales of the spectrum.
+The resulting forcing is built from a mixture of $1/3$ compressive and $2/3$ solenoidal modes, whose phases evolve following a stochastic Ornstein-Uhlenbeck process (Uhlenbeck and Ornstein, 1930; Federrath et al., 2010).
+Simulations start from homogenous ICs and are driven for $t=10$ (in units of sound-crossing time, $t_{sc}=L/c=1$), until a statistically stationary state with RMS Mach number $\mathcal{M}\sim 0.34$ is reached.
+3 Results
+We present the main results of the simulations carried out with SPH-EXA.
+Technical details, such as the integral approach to derivatives (IA) used to evaluate gradients, the use of generalized volume elements (GVE), the treatment of dissipation with both variable AV coefficients (VAV), and slope reconstruction technique (SLR), are deferred to the appendices.
+The appendices also include additional diagnostics based on the second-order velocity structure function, highlighting the advantages of SPH’s fully Lagrangian description.
+3.1 Impact of solver elements at fixed resolution
+To evaluate the weight and significance of the different elements that contribute to our SPH implementation, we conducted identical simulations while progressively incorporating more of these elements.
+Figure 1 shows the power spectrum of the velocity field of simulations of $250^{3}$ particles with increasing complexity, compared to the S3 simulation of (Bauer and Springel, 2012) with $256^{3}$ particles and the classic SPH formulation (i.e.
+gradients calculated directly as kernel derivatives, no GVEs, and no viscosity switches), labeled STD in our plot.
+Figure 3: Slice of thickness $2h$ centered on $z=0$ showing the velocity magnitude with increasing resolution (SPH-EXA). We represent particles directly.
+The inclusion of IA recovers the correct slope in the inertial range of the Kolmogorov cascade at low $k$, considerably extending it compared to STD (Valdarnini, 2016), but it does not remove the overly steep slope $\sim-4$ at large $k$.
+At these small scales something else is needed.
+Adding GVE and especially VAV moves in the right direction, revealing the bottleneck effect as an energy excess near the resolution limit, as commonly seen in mesh codes.
+Finally, the slope-limited reconstruction (SLR) provides the decisive improvement: it extends the inertial range almost a decade relative to STD, makes the bottleneck clearly visible, and yields a dissipative-range slope comparable to non-Lagrangian codes (see Fig.
+5).
+3.2 Resolution scaling and flow morphology
+In Fig.
+2 we present the power spectrum of the velocity field for simulations performed with SPH-EXA and with different particle counts at the end of each simulation ($t=10$).
+It is clear that with increasing resolution, the inertial range over the Kolmogorov scale is extended to larger wavenumber, and more importantly, that the rate of enhancement with increasing number of particles is not marginal or small, but considerable and consistent with that obtained with modern mesh-based and moving-mesh methods.
+This is a novel result, which means that SPH can indeed resolve smaller scales with higher resolution, in opposition to the results found with older SPH implementations (e.g., Bauer and Springel, 2012).
+Figure 3 shows a thin slice of the velocity field for increasing resolution with SPH-EXA, from $200^{3}$ up to $3200^{3}$, where it is evident how we can resolve a delicate and rich mixture of large- and small-scale features.
+Figure 4 compares the mass-weighted and volume-weighted density PDFs, with a fitted normal distribution.
+We can see the expected tail bias towards low densities, since in subsonic flows, local pressure increases in colliding flows can prevent the buildup of high-density regions (Konstandin et al., 2012).
+Figure 4: Instantaneous density PDF for the $400^{3}$ simulation at $t=10$. Comparison of mass-weighted and volume-weighted PDFs.
+3.3 Cross-code comparison
+We can now compare our results directly to other state-of-the-art codes: moving-mesh and meshless finite-volume methods.
+Figure 5 shows velocity power spectra from AREPO, GIZMO, and SPH-EXA.
+All three reproduce a comparable inertial range and a bottleneck of similar magnitude at comparable $k$.
+This suggests that, once the injection method and effective viscosity are matched, the cascade between the forcing and the dissipation wavenumbers behaves similarly across methods.
+Note that we compare runs with a similar number of fluid elements, which does not necessarily imply the same spatial resolution.
+In SPH, resolution is set by the smoothing length, which covers significantly more than just one particle.
+At $256^{3}$ cells, a proxy is the cell size, which for AREPO is $\sim 1/256\sim 3.9\times 10^{-3}$, comparable to our $400^{3}$ SPH simulation, with an average of $h\sim 3.6\times 10^{-3}$.
+Spatial resolution matters the most in the dissipation regime, where AREPO, GIZMO, and SPH-EXA have different transfer functions.
+In particular, GIZMO and SPH-EXA show a similar gentler slope in the roll-off, while AREPO is slightly steeper, pointing to a stronger high-$k$ damping.
+However, these differences are modest and the qualitative behavior persists at higher resolution, with the break shifting to larger $k$, as expected.
+Figure 5: Comparison of the velocity power spectra for different codes at $t=10$. AREPO and GIZMO curves are from their respective papers.
+Figure 6 shows, from top to bottom, the module of the velocity, density, and enstrophy of the fluid at the end of our simulations.
+The two columns on the left are an excerpt of the results with AREPO and the SPH used in (Bauer and Springel, 2012).
+It is clear from these two columns that the SPH used in that work was unable to resolve subsonic turbulence because it lacks all small-scale features.
+The column on the right shows the results of SPH-EXA with $250^{3}$ particles.
+The similarities with the results of AREPO are evident, and there is a rich mix of fine structures.
+This can also be seen in Fig.
+3, which illustrates a slice of the module of the velocity field at $t=10$ for different resolutions.
+SPH-EXA also produces an excellent second-order velocity structure function (appendix C).
+Additional colormaps of the highest resolution simulation, $3200^{3}$, can be found in appendix D.
+Figure 6: Comparison of SPH-EXA results with AREPO and SPH-tav in Bauer and Springel (2012) at $t=25$.
+From top to bottom: module of velocity, density, and enstrophy.
+The two columns on the left are an excerpt of the original results of AREPO and their SPH with reduced artificial viscosity (both with $256^{3}$ fluid elements) of Bauer and Springel (2012) shown in their Fig.
+4.
+The column on the right is the result of SPH-EXA with $250^{3}$ particles.
+The fields show directly particles in a thin slice around $Z=0$.
+No rasterization was performed.
+4 Summary and conclusions
+SPH-EXA combines a modern SPH formulation with state-of-the-art parallelization and GPU acceleration.
+This enables simulations that mitigate historical SPH limitations, such as excessive dissipation and inaccurate gradient evaluations, that have long hindered SPH from matching the performance of grid-based and moving-mesh codes in this field.
+Our findings not only align well with results from non-SPH codes like AREPO and GIZMO, but also illustrate the robustness of SPH in handling complex turbulent flows at resolutions previously thought unattainable with this method.
+We demonstrate that, with increasing resolution, the inertial range of the power spectra extends to larger wavenumbers, exhibiting features comparable to those seen in non-SPH codes.
+These improvements allow SPH to better exploit its intrinsic strengths: truly Lagrangian sampling (particles are the fluid elements, avoiding tracer advection and particle-to-grid interpolation noise), excellent conservation properties (exact mass advection and pairwise-conservative exchanges of momentum and energy), and flexibility across physical regimes.
+In particular, SPH couples naturally and accurately to tree methods for self-gravity, supporting its use in more demanding astrophysical contexts.
+Acknowledgements.
+This work has been supported by the Swiss Platform for Advanced Scientific Computing (PASC) project ”SPH-EXA: Optimizing Smoothed Particle Hydrodynamics for Exascale Computing”.
+It has also been carried out as part of the SKACH consortium through funding from SERI.
+We also acknowledge the financial support of the European Research Council through the ERC Synergy Grant “ECOGAL” (project ID 855130), from the German Excellence Strategy via the Heidelberg Cluster of Excellence (EXC 2181 - 390900948) “STRUCTURES”, and from the German Ministry for Economic Affairs and Climate Action in project “MAINN” (funding ID 50OO2206).
+The numerical calculations have been supported by the EuroHPC-JU Extreme Scale Access Mode through the project ”TGSF: The role of turbulence and gravity in star formation” (EHPC-EXT-2023E01-031).
+The authors acknowledge the support of the LUMI Supercomputer Center, the Swiss National Supercomputing Center (CSCS - allocations c32 and ch18), and the Center for Scientific Computing (sciCORE) at the University of Basel.
+Appendix A SPH Ingredients
+In this section, we discuss the ingredients of our SPH formulation that have the largest impact on our simulations of subsonic turbulence.
+Namely, the calculation of gradients, dissipation, and generalized volume elements.
+A.1 Gradients
+SPH-EXA implements an integral approach to calculate derivatives (García-Senz et al. 2012; Cabezón et al. 2012).
 The IA formalism calculates gradients of an arbitrary function $f$ by solving the following matrix equation:
 $\left[\begin{array}[]{c}\partial f/\partial x_{1}\\ \partial f/\partial x_{2}\\ \partial f/\partial x_{3}\\ \end{array}\right]_{a}=\left[\begin{array}[]{ccc}\tau_{11}&\tau_{12}&\tau_{13}\\ \tau_{21}&\tau_{22}&\tau_{23}\\ \tau_{31}&\tau_{32}&\tau_{33}\end{array}\right]^{-1}_{a}\left[\begin{array}[]{c}I_{1}\\ I_{2}\\ I_{3}\\ \end{array}\right]_{a}\,$, (1)
-where the $\tau$ tensor elements are defined as:
-$\tau_{ij,a}=\sum_{b}V_{b}~{}(x_{i,b}-x_{i,a})(x_{j,b}-x_{j,a})~{}W_{ab}(h_{a});~{}i,j=1,3\,.$ (2)
+where the $\tau$ tensor elements are defined as the weighted sum of the dyadic product of the relative distances $(\mathbf{x}_{b}-\mathbf{x}_{a})\otimes(\mathbf{x}_{b}-\mathbf{x}_{a})$:
+$\tau_{ij,a}=\sum_{b}V_{b}\penalty 10000\ (x_{i,b}-x_{i,a})(x_{j,b}-x_{j,a})\penalty 10000\ W_{ab}(h_{a});\penalty 10000\ i,j=1,3\,.$ (2)
 Here $x_{i}$ is the vector position coordinates, $V_{b}$ the volume element, $W_{ab}$ the interpolation kernel, and $I_{i,a}$ the vector integral elements, which are calculated as:
-$I_{i,a}=\left[\sum_{b}V_{b}f_{b}~{}(x_{i,b}-x_{i,a})~{}W_{ab}(h_{a})\right];~{}i=1,3\,.$ (3)
-As a consequence, the gradients calculation is:
-$\nabla_{i}f_{a}=\sum_{b}V_{b}(f_{b}-f_{a})A_{i,ab}(h_{a});~{}i=1,3\,$ (4)
+$I_{i,a}=\left[\sum_{b}V_{b}f_{b}\penalty 10000\ (x_{i,b}-x_{i,a})\penalty 10000\ W_{ab}(h_{a})\right];\penalty 10000\ i=1,3\,.$ (3)
+As a consequence, the gradient calculation is:
+$\nabla_{i}f_{a}=\sum_{b}V_{b}(f_{b}-f_{a})A_{i,ab}(h_{a});\penalty 10000\ i=1,3\,$, (4)
 where,
-$A_{i,ab}(h_{a})=\sum_{j=1}^{3}c_{ij,a}(h_{a})(x_{j,b}-x_{j,a})W_{ab}(h_{a})\,$ (5)
-where the summation is up to 3 for 3D simulations and $c_{ij,a}$ the elements of the inverse of matrix $\mathcal{T}$, that is, $\mathcal{C}=\mathcal{T}^{-1}$ (Eq. 1).
-II-B Artificial Viscosity
-The basic AV implementation in SPH-EXA is the following:
-$\Pi_{ab}^{AV}=\begin{cases}-\frac{\alpha}{2}v_{ab}^{sig}~{}w_{ab}&\text{for~{}~{}$\mathbf{x}_{ab}\cdot\mathbf{v}_{ab}<0$}\\ 0&\text{otherwise}\end{cases}\,,$ (6)
-where $\alpha$ is the viscous parameter and is usually set to 1. The signal velocity estimator between a pair of particles is:
+$A_{i,ab}(h_{a})=\sum_{j=1}^{3}c_{ij,a}(h_{a})(x_{j,b}-x_{j,a})W_{ab}(h_{a})\,$, (5)
+where the summation is up to 3 for 3D simulations and $c_{ij,a}$ the elements of the inverse of tensor $\mathcal{T}$, that is, $\mathcal{C}=\mathcal{T}^{-1}$ (Eq. 1).
+A.2 Dissipation
+The basic dissipation control in SPH-EXA is done using the following artificial viscosity (AV) prescription:
+$\Pi_{ab}^{AV}=\begin{cases}-\frac{\alpha}{2}v_{ab}^{sig}\penalty 10000\ w_{ab}&\text{for\penalty 10000\ \penalty 10000\ $\mathbf{x}_{ab}\cdot\mathbf{v}_{ab}<0$}\\ 0&\text{otherwise}\end{cases}\,$, (6)
+where $\alpha$ is the viscous parameter and is set to 1. The signal velocity estimator between a pair of particles is:
 $v^{sig}_{ab}=\frac{1}{2}(c_{a}+c_{b})-2w_{ab}\,$, (7)
@@
 We reconstruct the local velocity field for each particle by removing the linear component of the velocity jump that is used to trigger and calculate the AV in Eq.
-6 [19].
+6 (Frontiere et al.
+2017).
+This is the slope-limited reconstruction (SLR) method.
 Assuming that it is differentiable in the neighborhood of each SPH particle, the best linear approximation of the velocity field is given by multiplying its Jacobian by the distance to its neighbors.
@@
 where $\mathbf{x}^{T}_{ab}=\mathbf{x}_{a}-\mathbf{x}_{b}$ is the column vector of relative distance and $\mathbf{J}_{\mathbf{v}}$ is the Jacobian of the velocity field.
-In the case of the presence of discontinuities, such a linear estimator is not accurate enough; hence, the $\phi_{ab}$ factor is designed to go to zero in these cases, while it is kept close to 1 when the field is smooth.
-To this end, we use a van Leer-like limiter [19]:
+In the case of discontinuities, such a linear estimator is not accurate enough; hence, the $\phi_{ab}$ factor is designed to go to zero in these cases, while it is kept close to 1 when the field is smooth.
+To this end, we use a van Leer-like limiter (Frontiere et al.
+2017):
 $\phi_{ab}$ $=\max\left[0,\min\left(1,\frac{4F_{ab}}{(1+F_{ab})^{2}}\right)\right]\times\kappa_{ab}\,$, (10)
-$\kappa_{ab}$$\equiv\begin{dcases}\exp{\left[-\left(\frac{q_{ab}-q_{crit}}{q_{fold}}\right)^{2}\right]}&\text{if~{}~{}$q_{ab}<q_{crit}$}\,,\\ 1&\text{if~{}~{}$q_{ab}\geq q_{crit}$}\,,\end{dcases}$ (11)
+[…]
 $q_{ab}$ $\equiv\min(q_{a},q_{b})\,$, (12)
@@
 11) that is applied to the van Leer limiter (Eq.
-10) and we use the recommended value [19] of $q_{fold}=0.2$.
-The final step is to effectively control excessive dissipation.
+10) and we use the recommended value (Frontiere et al.
+2017) of $q_{fold}=0.2$.
+The final step is to control excessive dissipation effectively.
 This can be done by allowing a dynamic individual viscous parameter $\alpha_{a}$ in Eq.
 6, rather than having a constant value for all SPH particles.
-Several methods have been developed to calculate $\alpha_{a}$, but the nominal works of [23] and [24] are the inspiration behind our implementation in SPH-EXA.
+Several methods have been developed to calculate $\alpha_{a}$, but the nominal works of Cullen and Dehnen (2010) and Read and Hayfield (2012) are the inspiration behind our implementation in SPH-EXA.
 The main idea is that $\alpha_{a}$ should be initialized by default at a rather low value (usually $\alpha_{a}=0.05$).
 Then it increases rapidly if dissipation is needed and decays exponentially to its original value when it is not needed, effectively acting as a switch for the AV at the position of each SPH particle.
-We start by calculating a local viscous parameter $\alpha_{loc,a}$, following the definition [23]:
-$\alpha_{loc,a}=\begin{dcases}\frac{\alpha_{max}\mathcal{D}_{a}}{\mathcal{D}_{a}+h_{a}|\nabla\cdot\mathbf{v}_{a}|+0.05c_{a}}&\text{for~{}~{}$\nabla\cdot\mathbf{v}_{a}<0$}\,,\\ 0&\text{otherwise}\,,\end{dcases}$ (15)
+We start by calculating a local viscous parameter $\alpha_{loc,a}$, following the definition (Read and Hayfield 2012):
+[…]
 where $\mathcal{D}_{a}=h^{2}_{a}|\nabla(\nabla\cdot\mathbf{v}_{a})|$ predicts flow convergence by comparing the gradient of $\nabla\cdot\mathbf{v}_{a}$ with the local value of $\nabla\cdot\mathbf{v}_{a}$.
@@
   • if $\alpha_{loc,a}<\alpha_{a}$: we let $\alpha_{a}$ decay smoothly .
-In the case that $\alpha_{a}$ has to decay, we evaluate an instantaneous decay rate $\dot{\alpha}_{a}$ as in [23]:
-$\dot{\alpha}_{a}=\begin{dcases}(\alpha_{loc,a}-\alpha_{a})/\tau_{a}&\text{if~{}~{}$\alpha_{min}<\alpha_{loc,a}<\alpha_{a}$}\,,\\ (\alpha_{min}-\alpha_{a})/\tau_{a}&\text{if~{}~{}$\alpha_{min}\geq\alpha_{loc,a}$}\,.\end{dcases}$ (16)
+In the case that $\alpha_{a}$ has to decay, we evaluate an instantaneous decay rate $\dot{\alpha}_{a}$ as in (Read and Hayfield 2012):
+$\dot{\alpha}_{a}=\begin{dcases}(\alpha_{loc,a}-\alpha_{a})/\tau_{a}&\text{if\penalty 10000\ \penalty 10000\ $\alpha_{min}<\alpha_{loc,a}<\alpha_{a}$}\,,\\ (\alpha_{min}-\alpha_{a})/\tau_{a}&\text{if\penalty 10000\ \penalty 10000\ $\alpha_{min}\geq\alpha_{loc,a}$}\,.\end{dcases}$ (16)
 The decay time $\tau_{a}$ is defined as $\tau_{a}=h_{a}/v_{sig,max}$, and it is a function of the local spatial resolution and the maximum signal velocity $v_{sig,max}$ in the neighborhood of particle $a$, calculated according to Eq.
@@
 Also, there is no restriction for $\alpha_{min}$, which can be set to zero.
-II-C Volume elements
+In this respect, it is worth noticing the latest works pointing to the interesting possibility of not needing switches at all, if a modulated version of the SLR is used (Price 2024; Sandnes et al.
+2025; García-Senz and Cabezón 2025).
+This is achieved by taking constant $\alpha=1$ and multiplying the limiter $\phi_{ab}$ in Eqs.
+8 and 9 by:
+$1-\mathcal{B}^{p}\,$, (17)
+where $\mathcal{B}$ is the Balsara limiter,
+${\mathcal{B}}_{a}=\frac{|\nabla\cdot\mathbf{v}|_{a}}{|\nabla\cdot\mathbf{v}|_{a}+|\nabla\times\mathbf{v}|_{a}+10^{-4}\frac{c_{a}}{h_{a}}}\,.$ (18)
+García-Senz and Cabezón (2025) showed that taking $p=2$ in Eq.
+17 provides the best results across a series of tests involving shocks and/or shear, improving considerably over the use of switches and switches+SLR techniques.
+However, in order to get the best results in simulating subsonic turbulence, they recommend including an additional control in the amplitude of the AV to avoid spurious triggering of dissipation by the random motion of particles in the smaller scales.
+This can be done either with a switch, as shown before, or directly with a limiter, such as Balsara.
+A.3 Volume elements
 In the context of SPH, where we are discretizing the SPH integral interpolant, it becomes necessary to compute the volume element that corresponds to each particle.
 An uncomplicated yet highly effective method is $dx^{\prime 3}\equiv V_{b}\sim m_{b}/\rho_{b}$.
-However, it is known [25, 26] that the discretization of the volume element, present in most SPH equations, can be improved beyond this classical approximation using generalized volume elements (GVE).
+However, it is known (Saitoh and Makino 2013; Hopkins 2013) that the discretization of the volume element, present in most SPH equations, can be improved beyond this classical approximation using generalized volume elements (GVE).
 GVE are usually expressed as:
@@
 where $X$ is an arbitrary function.
-In SPH-EXA we adopted the GVE described in the work of [18], which has shown to provide a better treatment of discontinuities, stronger resistance to tensile instability, better partition of unity, and more accurate gradient evaluation:
+In SPH-EXA, we adopted the GVE described in the work of García-Senz et al.
+(2022), which has been shown to provide a better treatment of discontinuities, stronger resistance to tensile instability, better partition of unity, and more accurate gradient evaluation:
 $X_{a}=\frac{m_{a}}{\rho^{0}_{a}}\,$, (20)
-where $\rho^{0}_{a}=\sum_{b}m_{b}W_{ab}$ is the standard SPH density calculation. With Eqs. 17 and 18 we can simply calculate the particle density as $\rho_{a}=m_{a}/V_{a}$.
-II-D Lagrangian compatibility
-One of the most common artifacts in SPH simulations is tensile instability, which can, in some cases, prevent mixing and the development of hydrodynamical instabilities.
-Several approaches have been developed to deal with such problems [27, 28, 25], all of which require a departure to some extent from the pure Lagrangian formulation.
-In SPH-EXA we adopted a novel implementation [18], which, to our knowledge, is currently the only adaptive method that sacrifices Lagrangian compatibility exclusively in regions of the fluid at (and adjacent to) contact interphases, improving the accuracy at those regions, in turn.
-Hence, we define a self-adaptive parameter that is individual to each particle-neighbor pair as follows:
-$\sigma_{ab}\equiv\begin{cases}0&\text{if~{}~{}$At_{ab}\leq At_{min}$}\,,\\ R(At_{ab}-At_{min})&\text{if~{}~{}$At_{min}\leq At_{ab}\leq At_{max}$}\,,\\ 1&\text{if~{}~{}$At_{ab}\geq At_{max}$}\,,\end{cases}$ (19)
-with $R=(At_{max}-At_{min})^{-1}$ and $At_{ab}$ is the pairwise Atwood number, defined as $At_{ab}=\left|\frac{\rho_{a}-\rho_{b}}{\rho_{a}+\rho_{b}}\right|$.
-We use the values $At_{max}=0.2$ and $At_{min}=0.1$, as suggested in [18].
-In this implementation $0\leq\sigma\leq 1$ (see Eqs.
-22 and 23).
-In particular, when $\sigma=0$ the resulting SPH equations are fully Lagrangian, while $\sigma=1$ leads to a version less Lagrangian compatible but more tensile resistant.
-The parameter $\sigma$ appears in the momentum and energy SPH equations and has the effect of crossing and uncrossing the particle and neighbor’s magnitudes.
-II-E Interpolation kernel
-We use the sinc family ($S_{n}$) [29, 30] as interpolation kernels,
-$W_{ab}=S_{n}=\frac{K}{h_{a}^{d}}\begin{dcases}1&q_{a}=0\\ \left[\operatorname{sinc}\left(\frac{\pi}{2}q_{a}\right)\right]^{n}&q_{a}\leq 2\\ 0&q_{a}>2\\ \end{dcases}$ (20)
-where $K$ is a normalization constant, $h$ is the smoothing length, $d$ is the spatial dimension, $q_{a}$ is the normalized distance between neighbor particles ($q_{a}=|\mathbf{x}_{b}-\mathbf{x}_{a}|/h_{a}$), and $n$ is a user-defined real value that makes the kernel more centrally peaked as it increases.
-Usually $93<n<$ 9.
-It is known [31] that a necessary condition for an SPH kernel to avoid pairing instability is that its Fourier transform is definite positive.
-As such, they proposed the use of Wendland functions [32] as SPH kernels.
-These have been adopted by many researchers for this reason.
-The other side of the coin is that Wendland kernels need a large number of neighbors to perform accurate interpolations, which leads to considerable computational efforts.
-A possible alternative is to use pairing-resistant kernels.
-Despite not having an always positive Fourier transform, they are constructed to have small negative Fourier regions and push them to large wavenumbers.
-This allows these kernels to use a lower number of neighbors than Wendland functions, while still suppressing pairing instability.
-Sinc kernels are an example of such kernels.
-The simulations reported in this work were calculated choosing $n=6$ in Eq.
-20.
-II-F SPH-EXA hydrodynamic equations
+where $\rho^{0}_{a}=\sum_{b}m_{b}W_{ab}$ is the standard SPH density calculation. With Eqs. 19 and 20 we can simply calculate the particle density as $\rho_{a}=m_{a}/V_{a}$.
+A.4 SPH-EXA hydrodynamic equations
 The final set of hydrodynamic equations in SPH-EXA is as follows:
 $\rho_{a}=$ $m_{a}/V_{a}\,$, (21)
-$\begin{split}\frac{dv_{i,a}}{dt}=&-\sum_{b}m_{b}\times\\ &\left[\frac{X_{a}^{2-\sigma}X_{b}^{\sigma}P_{a}}{\Omega_{a}m_{a}^{2}~{}k_{a}}A_{i,ab}(h_{a})+\frac{X_{b}^{2-\sigma}X_{a}^{\sigma}P_{b}}{\Omega_{b}m_{b}^{2}~{}k_{b}}A_{i,ab}(h_{b})\right]\,,\end{split}$ (22)
-$\begin{split}\frac{du_{a}}{dt}=&\frac{X_{a}^{2-\sigma}P_{a}}{m_{a}^{2}\Omega_{a}k_{a}}\sum_{b}\sum^{d}_{i=1}m_{b}X_{b}^{\sigma}\left[(\mathbf{v}_{a}-\mathbf{v}_{b})\cdot\mathbf{A}_{ab}(h_{a})\right]\\ &+\left(\frac{du_{a}}{dt}\right)^{AV}\,.\end{split}$ (23)
+[…]
+$\begin{split}\frac{du_{a}}{dt}=&\frac{X_{a}^{2-\sigma}P_{a}}{m_{a}^{2}\Omega_{a}k_{a}}\sum_{b}\sum^{d}_{i=1}m_{b}X_{b}^{\sigma}\left[(\mathbf{v}_{a}-\mathbf{v}_{b})\cdot\mathbf{A}_{ab}(h_{a})\right]+\left(\frac{du_{a}}{dt}\right)^{AV}\,.\end{split}$ (23)
 Here $P_{a}$ is the pressure, $X_{a}=m_{a}/\rho^{0}_{a}$ the selected function for the GVE, $k_{a}$ the normalization of the GVE ($k_{a}\equiv\sum_{b}X_{b}W_{ab}$ in Eq.
-17), $A_{i,ab}(h_{a})$ the IA term, and $\Omega_{a}$ the grad-h term.
-The parameter $\sigma$ controls how close the implementation is to a pure Lagrangian scheme (as discussed above).
+19), $A_{i,ab}(h_{a})$ the IA term, and $\Omega_{a}$ the grad-h term.
+The parameter $\sigma$ controls how close the implementation is to a pure Lagrangian scheme (see García-Senz et al.
+2022).
 The last term in the energy equation includes the contributions of artificial viscosity.
-II-G Initial conditions
-We have designed an embedded initial conditions (IC) generator in SPH-EXA.
-In all our tests, we use a template of a relaxed box with $50^{3}$ particles.
-This box has previously been relaxed to achieve a glass-like configuration [33] with constant density and periodic boundary conditions.
-Replicating this box in different directions allows us to create on-the-fly initial conditions with a user-defined number of particles and in a distributed manner, only limited by the memory of the system.
-In this way, ICs can be scaled up just by changing one input argument and without the need to store large files, ensuring reproducibility and simplifying the process of performing suites of tests at scale.
-All simulations presented here were performed on a box of side $L=1$, periodic boundary conditions, initial density $\rho=1$, and ’quasi’-isothermal equation of state with an adiabatic index $\gamma=1.001$.
-The fluid is stirred using the Ornstein-Uhlenbeck process [34, 35] following a parabolic amplitude field $\sim 1-(k-k_{c})^{2}$, with $k=|\mathbf{k}|$, and $k_{c}=2$, which corresponds to a length $L/2$, and vanishing for $k\leq 1$ and $k\geq 3$.
-This approach is implemented to reduce the impact of the stirring on the energy cascade, facilitating a natural transfer of energy from larger to smaller scales.
-We also use a natural mixing of $1/3$ compressive and $2/3$ solenoidal contributions.
-All simulations start in a homogenous state and are stirred for a total of $t=10$, until they reach the stationary state [^1] All our times are normalized with the sound-crossing time, which in our setup is $t_{sc}=L/c=1$.
-with an RMS Mach number of $M=0.3$.
-III Results
-To evaluate the weight and significance of the different elements that contribute to our SPH implementation, we conducted identical simulations while progressively incorporating more of these elements.
-Figure 1 shows the power spectrum of the velocity field of simulations of $250^{3}$ particles with increasing complexity, compared to the S3 simulation of [9] with $256^{3}$ particles and the classic SPH formulation (i.e.
-gradients calculated directly as kernel derivatives, no GVEs, and no viscosity switches), labeled STD in our plot.
-Figure 1: Comparison of the power spectra of the velocity field for $M=0.3$ with different elements incorporated into the SPH solver.
-The dashed line is the theoretical Kolmogorov power spectrum of $\sim k^{-5/3}$.
-STD stands for standard SPH and is taken from [9] (simulation S3 in their Fig.
-6, with $256^{3}$ particles).
-All other simulations were performed with SPH-EXA and $250^{3}$ particles.
-IA stands for including the integral formalism to calculate derivatives.
-VAV stands for variable AV and implies the inclusion of AV switches.
-GVE is for generalized volume elements, and AVC for AV cleaning.
-All these elements have been described in Sec.
-II.
-The inclusion of IA recovers the correct slope in the inertial range of the Kolmogorov cascade for smaller wavenumbers, considerably extending it compared to STD, but is not enough to affect the classical excessive slope of $\sim-4$ at large wavenumbers.
-At these small scales something else is needed.
-The inclusion of GVE and artificial viscosity switches (labeled as VAV in Fig.
-1 provides a step in the right direction.
-With these elements, we start to see the bottleneck effect as the excess of energy at increasing wavenumber, commonly encountered close to the resolution limit in mesh-based simulations.
-Finally, the inclusion of the artificial viscosity cleaner (AVC in Fig.
-1) provides the final push, extending the inertial range almost a decade with respect STD, showing the bottleneck effect clearly, and having a slope in the dissipative regime similar to those of non-Lagrangian codes (see Fig.
-4).
-Figure 2 presents the power spectrum of the velocity field for simulations performed with SPH-EXA and with different particle counts at the end of each simulation ($t=10$).
-It is clear that with increasing resolution, the inertial range over the Kolmogorov scale is extended to larger wavenumber, and more importantly, that the rate of enhancement with increasing number of particles is not marginal or small, but considerable and consistent with that obtained with modern mesh-based and moving-mesh methods.
-This is a novel result, which means that SPH can indeed resolve smaller scales with higher resolution, in opposition to the results found with older SPH implementations by [9].
-Figure 3 shows a thin slice of the velocity field for increasing resolution with SPH-EXA, from $200^{3}$ up to $3000^{3}$, where it is evident how we can resolve a delicate and rich mixture of large- and small-scale features.
-Figure 2: Comparison of the power spectra of the velocity field at $t=10$ for $M=0.3$ at increasing resolution.
-Each line corresponds to an increase of a factor 2 in resolution with respect to the previous curve.
-Figure 3: Slice of the module of the velocity field with increasing resolution and calculated with SPH-EXA.
-Now we can directly compare our results with other state-of-the-art codes that are not purely Lagrangian.
-Figure 4 displays a comparison of the power spectra calculated with AREPO, GIZMO, and SPH-EXA.
-As can be seen, SPH-EXA is able to reproduce the same inertial range.
-We can also see the bottleneck effect to a very similar extent in all three codes.
-As this artificial excess of energy happens at a very similar wavenumber for similar resolutions, it points to the fact that the transition from the inertial range to the dissipation range happens at the same scale for a given resolution and for all three codes.
-It is worth noting that in Fig.
-4 we compare codes with a similar amount of fluid elements, which does not necessarily correspond to same resolution.
-In SPH codes, the spatial resolution is defined by the smoothing length, which covers significantly more than just one particle.
-At $256^{3}$ cells, the spatial resolution of AREPO roughly corresponds to $1/256\sim 3.9\times 10^{-3}$, which is in fact similar to our $400^{3}$ SPH simulation, which has an average smoothing length of $h\sim 3.6\times 10^{-3}$.
-This means that if we compare at the same resolution, and according to the results of Fig.
-4, SPH-EXA captures a considerably larger inertial range of the Kolmogorov cascade than AREPO.
-Likewise, the simulation of AREPO with $512^{3}$ cells has an approximate resolution of $\sim 1.9\times 10^{-3}$, which corresponds to our $800^{3}$ simulation ($h\sim 1.8\times 10^{-3}$) represented with a dashed line in Fig.
-4.
-Figure 4: Comparison of the power spectra of the velocity field for $M=0.3$ for different codes. AREPO [9] and GIZMO [20] curves are from their respective papers.
-Figure 5 shows, from top to bottom, the module of the velocity, density, and enstrophy of the fluid at the end of our simulations.
-The two columns on the left are an excerpt of the results with AREPO and the SPH used in [9].
-It is clear from these two columns that the SPH used in that work was not able to resolve subsonic turbulence, since it is missing all small-scale features.
-The two columns on the right show the results of SPH-EXA with $250^{3}$ and $1000^{3}$ particles.
-The similarities with the results of AREPO are evident, and there is a rich mix of fine structures that becomes richer with increasing resolution.
-This can also be seen in Fig.
-3, which illustrates a slice of the module of the velocity field at $t=10$ for different resolutions.
-Figure 5: Comparison of SPH-EXA results with AREPO and SPH-tav in [9].
-From top to bottom: module of velocity, density, and enstrophy.
-The two columns on the left are an excerpt of the original results of AREPO and their SPH with reduced artificial viscosity (both with $256^{3}$ fluid elements) of [9] shown in their Fig.
-4.
-The two columns on the right are the results of SPH-EXA with $250^{3}$ and $1000^{3}$
-Finally, Fig. 6 is a detailed image of a slice of the module of velocity of the $3000^{3}$ simulation where we can appreciate the very small-scale features resolved with SPH-EXA.
-Figure 6: Slice of the module of the velocity field at $t=10$ for a simulation with $3000^{3}$ particles. Color follows the same scale as Fig. 3 but in grey.
-IV Summary and conclusions
-SPH has matured enough to be competitive in simulations of subsonic turbulence, obtaining results comparable to Eulerian codes.
-In particular, SPH-EXA combines a modern SPH formulation with the latest techniques in parallelization and the use of accelerators.
-This allowed us to overcome historical limitations of SPH, such as excessive dissipation and inaccurate gradient evaluations, which have traditionally prevented SPH from matching the performance of grid-based and moving-mesh codes in this field.
-Our findings not only align well with results from non-Lagrangian codes like AREPO and GIZMO, but also illustrate the robustness of SPH in handling complex turbulent flows at resolutions previously thought unattainable with this method.
-We demonstrated that incorporating various components that address problems at multiple simulation scales collectively achieves these results.
-Isolated enhancements in the SPH framework cannot tackle the subsonic turbulence; however, a synergy of these improvements can.
-In particular, our combination of an integral approach to derivatives, generalized volume elements, artificial viscosity switches, and artificial viscosity cleaner shows excellent results.
-Finally, we proved that the inertial range of the power spectra significantly spans over larger vawenumbers with increasing resolution, exhibiting the same features as other non-Lagrangian state-of-the-art codes.
-These improvements have allowed SPH to finally tap into its inherent advantages, such as its intrinsic Lagrangian nature, excellent conservation properties, and adaptability to varied physical scenarios, thus paving the way for its application in more demanding astrophysical contexts.
-Acknowledgment
-This work has been supported by the Swiss Platform for Advanced Scientific Computing (PASC) project ”SPH-EXA: Optimizing Smoothed Particle Hydrodynamics for Exascale Computing”.
-It has also been carried out as part of the SKACH consortium through funding from SERI and supported by the Spanish MINECO grant PID2020-117252GB-100 and by the AGAUR/Generalitat de Catalunya grant SGR-386/2021.
-We also acknowledge the financial support of the European Research Council through the ERC Synergy Grant “ECOGAL” (project ID 855130), from the German Excellence Strategy via the Heidelberg Cluster of Excellence (EXC 2181 - 390900948) “STRUCTURES”, and from the German Ministry for Economic Affairs and Climate Action in project “MAINN” (funding ID 50OO2206).
-The numerical calculations have been supported by the EuroHPC-JU Extreme Scale Access Mode through the project ”TGSF: The role of turbulence and gravity in star formation: unveiling the sonic scale with Smoothed Particle Hydrodynamics” (proposal ID: EHPC-EXT-2023E01-031).
-The authors acknowledge the support of the LUMI Supercomputer Center, the Swiss National Supercomputing Center (CSCS - allocation c32), and the Center for Scientific Computing (sciCORE) at the University of Basel, where these calculations were performed.
+Appendix B The relevance of accurate grad-h terms
+Figure 7: Effect of the accuracy in grad-h terms.
+The top row shows the value of the grad-h correction $\Omega$ for all particles in a $250^{3}$ simulation at $t=10$.
+Left, $\Omega$ is calculated using the density field directly.
+Right, using the differences in the density field.
+The red lines show the mean value of $\Omega$.
+The bottom row shows a thin slice of the resulting density field.
+Grad-h terms are an important addition to the SPH technique when using variable smoothing lengths ($h$) (Springel and Hernquist 2002; Monaghan 2002).
+It takes into account the spatial variation of $h$ as a corrective term in the momentum (Eq.
+22) and energy (Eq.
+23) equations:
+$\Omega_{a}=\left[1+\frac{h_{a}}{3\rho_{a}}\frac{\partial\rho_{a}}{\partial h_{a}}\right]^{-1}\,.$ (24)
+These corrections are a standard method to improve accuracy and consistency in particle interpolation, reducing numerical instability in variable-density simulations, and as such, they are included by default in most astrophysics SPH codes.
+During the subsonic simulations presented above, we noticed that the density field appeared ”granulated” only when grad-h corrections were used jointly with SLR dissipation (see Fig.
+7, left column).
+Initially, we suspected that the new addition (the SLR dissipation control) was the source of the problem, but it turned out that the problematic term was how the grad-h correction was calculated.
+After inspecting the values of this correction we noticed that the average was systematically below 1 and with a relatively large spread.
+For example for the $250^{3}$ simulation, $\left<\Omega\right>=0.98\pm 0.16$.
+This doesn’t seem much, but it is exactly the kind of ”small” inconsistency that becomes relevant in subsonic turbulence, in particular when the dissipation formulation is designed to stop damping that noise.
+In low-Mach turbulence, physical density fluctuations are weak, so any residual inconsistency in the pressure–density–$h$ coupling becomes visible. We can write the grad-h field as:
+$\Omega_{a}=\bar{\Omega}+\delta\Omega_{a}\,.$ (25)
+Then, expanding the multiplicative term in the SPH equations:
+$\frac{1}{\Omega_{a}}=\frac{1}{\bar{\Omega}}\left(1-\frac{\delta\Omega_{a}}{\bar{\Omega}}+\mathcal{O}(\delta\Omega_{a}^{2})\right)$ (26)
+The $1/\Omega_{a}$ multiplicative factor that scales the pressure forces has two error sources: a biased mean and a noisy scatter.
+The first is just a constant scaling affecting the effective compressibility and shifting the global balance slightly.
+It shows as a systematic drift: slightly different RMS Mach, different effective forcing response, and slightly different spectra cutoff location, but doesn’t introduce particle-scale structure by itself.
+The second is the actual source of the granularity, since it is a multiplicative noise field coupled to the pressure acceleration.
+Since $\delta\Omega_{a}$ varies on the inter-particle scale and correlates with local disorder, it acts like a high-$k$ forcing of the compressive jitter:
+$\mathbf{a}_{a}^{P}\approx\frac{1}{\bar{\Omega}}\mathbf{a}^{P}_{\Omega=1}-\frac{1}{\bar{\Omega}^{2}}\delta\Omega_{a}\mathbf{a}^{P}_{\Omega=1}+\dots$ (27)
+If these corrections are in fact a noisy field, we get a noisy acceleration and particle jitter.
+This leads to noisy density estimates, noisy smoothing length, and back to noisy grad-h corrections, creating a feedback loop that distorts the field completely.
+This is usually masked by baseline AV, but becomes apparent when SLR suppresses dissipation in smooth regions.
+According to Eq.
+24 we need to evaluate $\partial\rho_{a}/\partial h_{a}$ and, as far as we know, the literature always implements it as a direct derivative of the density field.
+This is well known to have considerable E0 errors, and a typical trick is to include a scalar in the derivative to calculate it using the differences of the field, rather than the field directly.
+This has been used extensively to yield built-in conservation in SPH and zero derivatives when the field is constant.
+For example:
+$\frac{\partial\rho}{\partial h}=\frac{1}{\phi}\left[\frac{\partial(\phi\rho)}{\partial h}-\rho\frac{\partial\phi}{\partial h}\right]\,$, (28)
+where $\phi$ could be any other field or scalar. If we simply choose $\phi=1$:
+$\frac{\partial\rho_{a}}{\partial h_{a}}=\sum_{b}V_{b}\rho_{b}\frac{\partial W_{ab}(h_{a})}{\partial h_{a}}-\rho_{a}\sum_{b}V_{b}\frac{\partial W_{ab}(h_{a})}{\partial h_{a}}\,.$ (29)
+The first term on the RHS of Eq.
+29 is the common grad-h term that is used to calculate $\Omega_{a}$.
+The second term in the RHS of Eq.
+29 is the additional correction to evaluate grad-h as differences and reduce its noise.
+The practical implementation of this term depends on which volume elements are used.
+In SPH-EXA, we corrected the calculation of $\partial\rho_{a}/\partial h_{a}$, amending formula A17 from García-Senz et al.
+(2022).
+Defining:
+$\Psi_{a}\equiv\rho_{a}\sum_{b}V_{b}\frac{\partial W_{ab}(h_{a})}{\partial h_{a}}\,$, (30)
+we obtain:
+$\begin{split}\frac{\partial\rho_{a}}{\partial h_{a}}=\left[\frac{\rho_{a}}{\rho^{0}_{a}}-X_{a}W_{aa}(h_{a})\right]&\sum_{b}m_{b}\frac{\partial W_{ab}(h_{a})}{\partial h_{a}}-\rho^{0}_{a}\Psi_{a}+\\ \frac{m_{a}}{X_{a}}&\sum_{b}X_{b}\frac{\partial W_{ab}(h_{a})}{\partial h_{a}}-\rho_{a}\Psi_{a}\,.\end{split}$ (31)
+Note that if this formulation is used, self-contributions of particle $a$ must not be included.
+There are different ways to implement this correction, as well as different possibilities for the field $\phi$.
+We opted for the easiest and least invasive methodology.
+As a result, Fig. 7 (right column) shows that the average value of $\Omega$ improved considerably, becoming $\left<\Omega\right>=1.00\pm 0.04$, and the density field was recovered.
+Appendix C Second-order velocity structure function
+Figure 8: Mass-weighted turbulent Mach number as a function of spatial scale for a turbulent box simulation with a resolution of $2000^{3}$ particles.
+The dotted line indicates the RMS velocity of the simulation.
+The dashed line is a power-law fit to the turbulent cascade region.
+The SPH method is particularly advantageous for a wide range of mass-weighted analysis methods.
+Most importantly, it removes the need for special tracer particles as the SPH particles themselves serve as tracers.
+This allows for an easy computation of several quantities like the finite-time Lyapunov exponent and mass-weighted structure functions.
+The structure function is defined as
+${SF}_{2}(d/L)=\left\langle|\@vec{v}(\@vec{r})-\@vec{v}(\@vec{r}+\@vec{d}/L)|^{2}\right\rangle_{\@vec{r}}$ (32)
+where d=|d(r)|d=|\@vec{d}(\@vec{r})| is the distance between the two particles in a pair, and we compute the average over many SPH particle pairs, which makes this a mass-weighted structure function.
+$L$ is the turbulent driving scale.
+From this, one can compute the turbulent Mach number as
+$M_{t}(d/L)=\sqrt{{SF}_{2}(d/L)/(2c_{s}^{2})}$ (33)
+As an example, Figure 8 shows the turbulent Mach number as a function of the spatial scale for one of our SPH-EXA simulations.
+The computation was done using $10^{9}$ particle pairs per snapshot and was averaged over $41$ snapshots of the simulation, which is sufficient to achieve statistical convergence.
+In the regime of the turbulent cascade, we find a power law with an exponent of $0.333$.
+This perfectly matches the expected result of $1/3$ for subsonic Kolmogorov turbulence.
+Together with the density PDF shown in Figure 4, this demonstrates that SPH-EXA is capable of producing accurate, mass-weighted statistics for subsonic turbulence in a much more natural way than comparable mesh codes.
+Appendix D High-resolution view of the velocity field
+Figures 9 to 11 show a detailed image of a slice of the velocity magnitude, density, and enstrophy for the $3200^{3}$ simulation, respectively, where we can appreciate the very small-scale features resolved with SPH-EXA.
+All slices have a thickness of $2h$ and they are centered on $z=0$.
+Particles are displayed directly (no particle-to-grid interpolation is applied) and each slice contains $\sim 5.84\times 10^{7}$ particles.
+Figure 9: Velocity magnitude $|\mathbf{v}|$ at $t=10$ for the $3200^{3}$ simulation, shown in a slice of thickness $2h$ centered on $z=0$.
+Figure 10: Density at $t=10$ for the $3200^{3}$ simulation, shown in a slice of thickness $2h$ centered on $z=0$.
+Figure 11: Enstrophy at $t=10$ for the $3200^{3}$ simulation, shown in a slice of thickness $2h$ centered on $z=0$.